### v1

# ChainRank-DPO: Chain Rank Direct Preference Optimization for LLM Rankers

Abstract
Large language models (LLMs) have demonstrated remarkable effectiveness in text reranking through works like RankGPT, leveraging their human-like reasoning about relevance. However, supervised fine-tuning for ranking often diminishes these models’ general-purpose capabilities, including the crucial reasoning abilities that make them valuable for ranking. We introduce a novel approach integrating Chain-of-Thought prompting with an SFT-DPO (Supervised Fine-Tuning followed by Direct Preference Optimization) pipeline to preserve these capabilities while improving ranking performance. Our experiments on TREC 2019 and 2020 Deep Learning datasets show that our approach outperforms the state-of-the-art RankZephyr while maintaining strong performance on the Massive Multitask Language Understanding (MMLU) benchmark, demonstrating effective preservation of general-purpose capabilities through thoughtful fine-tuning strategies. Our code and data will be publicly released upon the acceptance of the paper.
ChainRank-DPO: Chain Rank Direct Preference Optimization for
LLM Rankers
Haowei Liu Santa Clara University Santa Clara, CA hliu6@scu.edu                        Xuyang Wu Santa Clara University Santa Clara, CA xwu5@scu.edu                        Guohao Sun Rochester Institute of Technology Rochester, NY gs4288@rit.edu
Zhiqiang Tao Rochester Institute of Technology Rochester, NY zhiqiang.tao@rit.edu                        Yi Fang Santa Clara University Santa Clara, CA yfang@scu.edu
1 Introduction
Text reranking is a vital task in information retrieval Liu (2009); Hasanain (2018), crucial for search engines Li et al. (2024), conversational AI Becker et al. (2012), and recommendation systems Zhao and Liu (2024). Large language models (LLMs) excel in reranking due to their reasoning and human-like thinking capabilities, enabling them to handle complex queries and ambiguous contexts. RankGPT Achiam et al. (2023) has set a high standard in listwise reranking Ma et al. (2019), leveraging its interpretability and systematic reasoning Brown et al. (2020) to advance reranking technologies.
Figure 1: While both methods can rerank passages, general methods struggle with simple QA tasks, whereas ChainRank successfully solves them step by step.
Building on the success of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT). RankVicuna Pradeep et al. (2023a), for instance, uses distilled data generated exclusively from GPT-3.5, focusing on efficient fine-tuning to improve ranking performance while maintaining a lightweight architecture. Moreover, RankZephyr Pradeep et al. (2023b) employs a more comprehensive approach utilizing distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results. Although these models achieve state-of-the-art performance in specific benchmarks, they highlight a key challenge in ranking: the trade-off between optimizing task-specific performance and preserving the broader reasoning capabilities of LLMs. Addressing this challenge requires novel strategies to balance specialization with generalization, ensuring that ranking models retain their versatility across diverse tasks.
Our investigations in Figure 1 reveal that while RankVicuna and RankZephyr exhibit strong performance in listwise ranking tasks, they have forfeited their mathematical reasoning capabilities following fine-tuning. This is evident from the figure, where the models fail to produce coherent solutions to mathematical problems, often outputting unrelated or nonsensical numerical sequences. Moreover, further analysis uncovered that these models have also lost their general text-generating capabilities. Regardless of the input prompt, the models struggle to produce meaningful or contextually appropriate responses, highlighting a significant trade-off introduced by their specialized fine-tuning for ranking tasks. Furthermore, while they exhibit comprehension of listwise ranking, they encountered difficulties with pairwise ranking, as their pairwise outcomes did not correspond with those generated by the listwise ranking Peng et al. (2024).
To address this limitation, we propose ChainRank strategy, built on LLaMA3-8b-instruct Dubey et al. (2024). Using the same training data as RankZephyr and RankGPT3.5/RankGPT4 as teacher models, we introduced a chain-of-thought (CoT) prompt to guide sequential passage ranking. We then implemented ChainRank-DPO, enhancing the model’s reasoning abilities for superior ranking performance. While traditional Direct Preference Optimization (DPO) Rafailov et al. (2024) performs well on chat benchmarks, it struggles with long-chain reasoning tasks like math and ranking due to error propagation. Inspired by Step-DPO Lai et al. (2024), we designed a novel DPO framework using overlapping ranking orders as the reward function, allowing better error correction. ChainRank preserves general text generation and reasoning capabilities, overcoming limitations in RankVicuna and RankZephyr. It outperforms both on TREC 2019/2020 Deep Learning Tracks Craswell et al. (2021, 2020) and BEIR benchmarks Thakur et al. (2021). We summarize the contributions of this work as follows:
  • We propose a novel chain-of-thoughts instruction (reranking) tuning that enables LLMs to rank passages based on relevance step by step.
  • To our best knowledge, this is the first work to introduce the Supervised Fine-Tune–DPO pipeline in the context of text reranking.
  • Using the same training data, our model outperforms both RankZephyr and the teacher model, RankGPT4, while retaining the general capabilities of LLMs.
{mdframed}
[ linecolor=black!60, linewidth=1pt, roundcorner=10pt, backgroundcolor=gray!5, shadow=true, shadowsize=5pt, shadowcolor=black!40, skipabove=10pt, skipbelow=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt ] USER: I will provide you with {num} passages, each indicated by a numerical identifier []. Rank the passages based on their relevance to the search query: {query}.
[1] {passage 1}
[2] {passage 2}
[{num}] {passage {num}}
Search Query: {query}.
Rank the {num} passages by selecting the most relevant passage at each step from the remaining passages. After choosing the most relevant passage, remove it from the pool and continue ranking until all passages are ordered.
Instructions:
Start with the most relevant passage and select it from the full list.
For each following step, pick the most relevant passage from the remaining passages only.
List the selected passages by their identifiers at each step, one after the other, until all passages are ranked.
Example Output:
Step 1: [4]
Step 2: [4, 2]
Step 3: [4, 2, 3]
step {num}: [4,2,3,15,…,14]
Final Answer: [4, 2, 3,…, 14]
Only respond with each step and the final answer, ensuring each passage is included once and ranked in descending relevance.
Figure 2: ChainRank Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively. The prompt ensures step-by-step selection, removal, and ordering of passages, with an example illustrating the expected output format.
2 Methodology
Figure 3: Training framework of our model. $Left:$ Stage 1 aims to teach the student model to perform text reranking. $Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform Chain DPO training.
2.1 CoT Reranking Prompt Design
Recent advancements in listwise text reranking, such as RankVicuna Pradeep et al. (2023a), RankZephyr Pradeep et al. (2023b), and related methods Liu et al. (2024), build on zero-shot techniques Ma et al. (2023); Sun et al. (2023), using prompt templates to reorder documents for metrics like nDCG Järvelin and Kekäläinen (2002). Given a query $q$ and documents ${d_{1},d_{2},...,d_{n}}$, the goal is to return a reordered list (e.g., $[1]>[2]>[3]...>[20]$).
Our ChainRank strategy frames listwise reranking as a chain-of-thought (CoT) reasoning task, selecting the most relevant document iteratively until all are ranked. The CoT prompt for LLaMA is shown in Figure 2.
2.2 Training Dataset
We train ChainRank model using 35k GPT-3.5 and 5k GPT-4 labeled instances from Pradeep et al. (2023b), derived from randomly selected MS MARCO v1 queries. Pyserini Lin et al. (2021) retrieved 20 BM25 candidates per query, which RankGPT3.5 and RankGPT4 ordered as teacher models. Malformed generations, such as missing or duplicate identifiers, were removed to improve data quality. We randomly selected 90% of the query-document pairs for Stage 1 SFT training and 10% for Stage 2 DPO training, denoted as SFT data and DPO data.
2.3 Training Stage 1: Supervised Fine-Tune
As shown in Figure 3, during the Supervised Fine-Tuning (SFT) phase, we use SFT data, maintaining a zero-shot setup since RankGPT and ChainRank do not rely on human-labeled data.
Before fine-tuning, the original LLaMA3-8B-Instruct model Dubey et al. (2024) failed to produce meaningful CoT reranking, simply replicating the example format without relevance-based results, highlighting the need for fine-tuning. The model weights are publicly available on HuggingFace. [^1] https://huggingface.co/meta-LLaMA/Meta-LLaMA-3-8B-Instruct
We fully fine-tune the 8B parameter LLaMA-3 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format. Training on four NVIDIA A100 80GB GPUs took approximately 39 hours.
2.4 Training Stage 2: Chain DPO
After Stage 1, the ChainRank-SFT model generates three ranking predictions ($y$) on prompts ($x$) from DPO training data. Predictions are evaluated by overlapping ranking orders with ground-truth labels, creating a preference dataset with prompts (x), chosen steps ($s_{w}$), rejected steps ($s_{l}$), and overlapping steps ($s_{o}$). Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y={s_{1},s_{2},\dots,s_{n}}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$. Overlapping steps ($s_{o}$) are tracked until a divergence is found, after which only the initial contiguous overlaps are included in $s_{o}$, and later steps are categorized as chosen ($s_{w}$) or rejected ($s_{l}$). This results in a dataset of $(x,s_{w},s_{l},s_{o})$.
The objective maximizes the likelihood of correct steps ($s_{w}$) while minimizing incorrect ones ($s_{l}$), using the loss:
$\begin{split}&\mathcal{L}(\theta)=-\mathbb{E}_{(x,s_{w},s_{l},s_{o})\sim D}\Big{[}\log\sigma\Big{(}\\ &\beta\log\frac{\pi_{\theta}(s_{w}\mid x;s_{o})}{\pi_{ref}(s_{w}\mid x;s_{o})}-\beta\log\frac{\pi_{\theta}(s_{l}\mid x;s_{o})}{\pi_{ref}(s_{l}\mid x;s_{o})}\Big{)}\Big{]}\,,\end{split}$
where $\pi_{\theta}(s\mid x;s_{o})$ represents the model’s probabilities, optimized to favor correct steps while leveraging reference model probabilities ($\pi_{\text{ref}}$) for stable training and better generalization, especially in zero-shot scenarios. This improves CoT reasoning’s sequential decision-making and model consistency. The Chain DPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately eight hours.
3 Experiments
3.1 Research Questions
Before conducting the experiments, we formulate several research questions to ensure that ChainRank could effectively address the challenges:
-RQ1: Does the CoT instruction tuning improve the step-by-step relevance ranking of passages compared to traditional ranking methods?
-RQ2: How does our model’s performance on text reranking tasks compare with existing models like RankZephyr and RankGPT4 across various datasets and task settings?
-RQ3: Does introducing DPO improve the performance and robustness of our model?
3.2 Evaluation Benchmarks
We evaluate ranking capabilities using TREC DL19 Craswell et al. (2020) and DL20 Craswell et al. (2021) Tracks, derived from MS MARCO V1 Bajaj et al. (2016), with human-annotated relevance labels. TREC DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 candidate passages retrieved by BM25. Additionally, we include cross-domain datasets from BEIR Thakur et al. (2021), not seen by the fine-tuned LLMs, for broader analysis. Using a sliding window strategy with a size of 20 and a stride of 10, we rerank 100 passages per query.
We use the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2020) to evaluate whether our model retains its text understanding abilities after fine-tuning for ranking tasks. MMLU tests proficiency across 57 diverse subjects, from elementary to advanced professional topics. We compare our model’s performance on MMLU with baseline LLMs to ensure ranking fine-tuning does not compromise general-purpose language understanding and generation capabilities, maintaining its versatility in broader tasks.
3.3 Results
Table 1 compares model performance on TREC, BEIR, and MMLU benchmarks. Our 8B ChainRank model outperforms baselines, including RankGPT4  [^2] RankGPT is the only closed-source and largest model in the table., even with 90% of the training data, and surpasses RankLLaMA3 under identical training conditions, showcasing the effectiveness of our CoT reranking prompt. In the DPO stage, ChainRank-DPO improves across all datasets with just one training epoch, enhancing robustness to passage variations (Appendix B).
On MMLU, ChainRank matches LLaMA3’s performance, while RankLLaMA3 shows a slight drop and RankVicuna performs poorly. RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU. Additional examples are in Appendix D.
We present a figure to illustrate the trade-off between performance and inference cost when generating ranking orders at varying step intervals. Notably, our method achieves comparable latency to RankLLaMA3 (26.5 vs. 25.0 seconds). By utilizing parallel distributed evaluation, we further reduce inference latency to 6.73 seconds on 4 A100 GPUs, demonstrating the efficiency of our approach.
Models TREC BEIR MMLU DL19 DL20 NFC COVID FIQA AVG. RankGPT4 0.746 0.708 0.406 0.749 0.333 0.864 BM25 0.506 0.480 0.325 0.595 0.236 N/A Contriever 0.616 0.599 0.328 0.596 0.329 N/A Gemma-7B 0.533 0.530 0.338 0.573 0.236 0.649 LLaMA3 0.641 0.621 0.269 0.628 0.216 0.662 RankZephyr 0.742 0.709 0.331 0.592 0.230 0.000 RankVicuna 0.668 0.655 0.338 0.592 0.236 0.373 RankLLaMA3 0.730 0.635 0.335 0.672 0.330 0.628 ChainRank-SFT 0.752 0.714 0.353 0.766 0.342 0.663 ChainRank-DPO 0.755 0.717 0.358 0.772 0.342 0.663
Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks. All the reranking tasks are based on BM25 retrieval results. The bold values in the table highlight the best performance across the respective benchmarks.
Figure 4: FLOPs and nDCG@10 performance across different step intervals. The blue line represents the inference cost (FLOPs), while the orange line shows the ranking performance (NDCG@10). As the step interval increases, FLOPs decrease significantly, with a slight drop in nDCG@10 scores.
4 Conclusion and Future Work
In this paper, we propose a novel approach, ChainRank, a novel zero-shot listwise text reranking model built on LLaMA3. We demonstrated the effectiveness of CoT reranking prompt and our SFT–Chain DPO pipeline. The results of our experiments show that our model achieves superior performance compared to many open-source and closed-source LLMs, such as RankZephyr and RankGPT4.
In future work, we plan to explore the application of our method to other models, such as Mistral, Zephyr, and LLaMA3.1, to evaluate its generalizability and performance across different architectures. Additionally, our future research could incorporate higher-quality datasets with varying numbers of passages per instance to enhance the diversity and robustness of the training data. Expanding the dataset in this manner could lead to improved overall performance.
5 Limitations
While this study provides valuable insights, several limitations should be acknowledged. Firstly, the need to include long examples of CoT formatting (Step1: …) in our prompts leads to increased training and inference times.
Secondly, although LLaMA3 has a larger context window (8192 tokens) compared to Vicuna and Zephyr (4096 tokens), we are still limited by the maximum number of passages in our training data, which is set to 20. As a result, we are restricted to ranking within a window size of 20 and require nine passes to complete the reranking of 100 passages.
Appendix A Baseline Selection
To illustrate the efficacy of ChainRank, we select several baseline models, including the existing retriever models, BM25 Robertson and Zaragoza (2009) and Contriever Izacard et al. (2021), zero-shot prompt-based LLMs, RankGPT4 and Gemma-7B Team et al. (2024), fine-tuned LLM-based models (RankVicuna, RankZephyr), as well as our base model LLaMA3-8B-Instruct. To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankVicuna but use our base model LLaMA3-8B-Instruct, denoted as RankLLaMA3.
Appendix B Window Size
Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 RankLLaMA3 0.567 0.663 0.666 ChainRank-SFT 0.697 0.737 0.755 ChainRank-DPO 0.671 0.755 0.758
Table 2: Performance of different models with different window size and step size, BM25 performs all of the retrieval stage, and all of the evaluations are performed on DL19
Appendix C Explanation of CoT strategies
To evaluate the trade-off between performance and inference cost, we trained our model to generate ranking orders after every 1, 3, 5, and 7 steps. An additional instruction was introduced to guide the model in returning ranking orders at these specified intervals. For comparison, we also included the performance of RankLLaMA3 to highlight the effectiveness of our CoT strategies
Appendix D Examples of general LLM tasks
To illustrate the degradation in general text understanding and generation capabilities of RankVicuna and Rankzephyr, we provide two examples below. Due to the length of ChainRank’s response, we have omitted the middle portion of the answer.
D.1 Example 1
Prompt Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions.
RankVicuna Ultimately, showcasing Hawaii’s rich history and cultural significance. Title: Embracing Hawaiian Culture: Discovering Historical Landmarks and Rich Traditions on the Big Island
RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [11] > [12] > [13]
ChainRank Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me. I had dreamed of visiting Hawaii for years, and finally, I was here. The warm tropical air enveloped me, and the sweet scent of plumeria filled my senses as I made my way to my hotel. Little did I know, this trip would be a journey not only to breathtaking natural wonders but also to a rich cultural heritage that would leave me feeling grateful and inspired. Immersing in the Local Culture: (Content omitted for brevity.) Must-See Attractions: (Content omitted for brevity.) A Taste of Aloha: (Content omitted for brevity.) Conclusion: (Content omitted for brevity.) Practical Tips and Recommendations: Visit Pearl Harbor and USS Arizona Memorial in the morning to avoid the crowds. Try local specialties like poke, laulau, and shave ice. Take a surf lesson or snorkeling trip to explore the islands’ stunning coastlines. Visit the Bishop Museum to learn about the history and geology of Hawaii. Explore the islands’ many farmers’ markets for fresh produce and local products. Learn traditional Hawaiian crafts, like leis and ukulele strings, from local artisans. Mahalo (thank you) for joining me on this journey to the Aloha State. I hope my story has inspired you to plan your own trip to Hawaii and experience the warmth and beauty of this incredible destination.
D.2 Example 2
Prompt Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided?
RankVicuna (Nothing was generated from the model.)
RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [3] > [4] > [5] > [6]
ChainRank I’d be happy to help you write a resignation letter to your current employer. Here’s a sample template you can use as a guide: Dear [Recipient’s Name], I am writing to inform you of my decision to resign from my position as [Your Position] at [Company Name], effective [Date of Last Day of Work]. I want to express my deepest gratitude to you and the entire [Company Name] team for the opportunities I’ve had during my time here. (Content omitted for brevity.) Please don’t hesitate to reach out to me if you need anything after my departure. I can be reached at [Your Email Address] or [Your Phone Number]. Sincerely, [Your Signature] Some tips to keep in mind when writing your resignation letter: Be professional and respectful: Use a formal tone and language throughout the letter. Express gratitude: Let your employer know how much you appreciate the opportunities you’ve had and the experience you’ve gained. Offer to help with the transition: Show that you’re committed to making the transition as smooth as possible by offering to complete outstanding tasks and assist with the handover of your responsibilities. Keep it brief: Aim for a letter that is no more than one page in length. Proofread: Make sure to proofread your letter carefully to ensure it is free of errors and typos. Remember to customize the letter to fit your own circumstances and experiences. Good luck with your resignation, and I hope it leads to a positive and successful transition for you!

### v2

@@
-ChainRank-DPO: Chain Rank Direct Preference Optimization for LLM Rankers
+RaCT: Ranking-aware Chain-of-Thought Optimization for LLMs
 Abstract
-Large language models (LLMs) have demonstrated remarkable effectiveness in text reranking through works like RankGPT, leveraging their human-like reasoning about relevance.
-However, supervised fine-tuning for ranking often diminishes these models’ general-purpose capabilities, including the crucial reasoning abilities that make them valuable for ranking.
-We introduce a novel approach integrating Chain-of-Thought prompting with an SFT-DPO (Supervised Fine-Tuning followed by Direct Preference Optimization) pipeline to preserve these capabilities while improving ranking performance.
-Our experiments on TREC 2019 and 2020 Deep Learning datasets show that our approach outperforms the state-of-the-art RankZephyr while maintaining strong performance on the Massive Multitask Language Understanding (MMLU) benchmark, demonstrating effective preservation of general-purpose capabilities through thoughtful fine-tuning strategies.
-Our code and data will be publicly released upon the acceptance of the paper.
-ChainRank-DPO: Chain Rank Direct Preference Optimization for
-LLM Rankers
+Large language models (LLMs) have shown significant promise in text reranking tasks by leveraging their advanced language understanding and reasoning capabilities.
+However, traditional supervised fine-tuning (SFT) approaches by ranking utilities can compromise LLMs’ general-purpose abilities.
+To address this challenge, we propose a novel LLM-based reranking algorithm – RaCT – that implements SFT with Chain-of-Thought prompting, followed by a ranking preference optimization (RPO).
+The proposed RaCT aims to enhance ranking performance for LLMs while preserving their inherent language modeling abilities.
+Experimental evaluations on the three public ranking benchmarks (TREC DL, BEIR, and BRIGHT) and one LLM benchmark demonstrate the superior ranking performance of RaCT with a retained language understanding and reasoning capacity.
+RaCT: Ranking-aware Chain-of-Thought Optimization for LLMs
 Haowei Liu Santa Clara University Santa Clara, CA hliu6@scu.edu                        Xuyang Wu Santa Clara University Santa Clara, CA xwu5@scu.edu                        Guohao Sun Rochester Institute of Technology Rochester, NY gs4288@rit.edu
@@
 (2023) has set a high standard in listwise reranking Ma et al.
-(2019), leveraging its interpretability and systematic reasoning Brown et al.
+(2019), leveraging its systematic reasoning Brown et al.
 (2020) to advance reranking technologies.
-Figure 1: While both methods can rerank passages, general methods struggle with simple QA tasks, whereas ChainRank successfully solves them step by step.
-Building on the success of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT).
-RankVicuna Pradeep et al.
-(2023a), for instance, uses distilled data generated exclusively from GPT-3.5, focusing on efficient fine-tuning to improve ranking performance while maintaining a lightweight architecture.
-Moreover, RankZephyr Pradeep et al.
-(2023b) employs a more comprehensive approach utilizing distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results.
-Although these models achieve state-of-the-art performance in specific benchmarks, they highlight a key challenge in ranking: the trade-off between optimizing task-specific performance and preserving the broader reasoning capabilities of LLMs.
-Addressing this challenge requires novel strategies to balance specialization with generalization, ensuring that ranking models retain their versatility across diverse tasks.
-Our investigations in Figure 1 reveal that while RankVicuna and RankZephyr exhibit strong performance in listwise ranking tasks, they have forfeited their mathematical reasoning capabilities following fine-tuning.
-This is evident from the figure, where the models fail to produce coherent solutions to mathematical problems, often outputting unrelated or nonsensical numerical sequences.
-Moreover, further analysis uncovered that these models have also lost their general text-generating capabilities.
-Regardless of the input prompt, the models struggle to produce meaningful or contextually appropriate responses, highlighting a significant trade-off introduced by their specialized fine-tuning for ranking tasks.
-Furthermore, while they exhibit comprehension of listwise ranking, they encountered difficulties with pairwise ranking, as their pairwise outcomes did not correspond with those generated by the listwise ranking Peng et al.
-(2024).
-To address this limitation, we propose ChainRank strategy, built on LLaMA3-8b-instruct Dubey et al.
-(2024).
-Using the same training data as RankZephyr and RankGPT3.5/RankGPT4 as teacher models, we introduced a chain-of-thought (CoT) prompt to guide sequential passage ranking.
-We then implemented ChainRank-DPO, enhancing the model’s reasoning abilities for superior ranking performance.
-While traditional Direct Preference Optimization (DPO) Rafailov et al.
-(2024) performs well on chat benchmarks, it struggles with long-chain reasoning tasks like math and ranking due to error propagation.
-Inspired by Step-DPO Lai et al.
-(2024), we designed a novel DPO framework using overlapping ranking orders as the reward function, allowing better error correction.
-ChainRank preserves general text generation and reasoning capabilities, overcoming limitations in RankVicuna and RankZephyr.
-It outperforms both on TREC 2019/2020 Deep Learning Tracks Craswell et al.
-(2021, 2020) and BEIR benchmarks Thakur et al.
-(2021).
+Figure 1: General text reranking LLMs excel at ranking tasks but struggle with general tasks such as open-ended text generation.
+General LLMs (e.g., LLaMA) perform well on diverse tasks but lack strong ranking capabilities.
+RaCT effectively integrates both, achieving high performance in ranking while maintaining general language generation abilities.
+Building on top of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT).
+To name a few, RankVicuna Pradeep et al.
+(2023a) uses distilled data generated exclusively from GPT-3.5 and focuses on efficient fine-tuning to improve ranking performance with a lightweight architecture.
+RankZephyr Pradeep et al.
+(2023b) develops a more comprehensive approach that utilizes distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results.
+While these models exhibit state-of-the-art ranking performance, they pose a new challenge for LLM rerankers: the trade-off between optimizing ranking utility and preserving the inherent language modeling capabilities.
+Empirically, over-emphasizing ranking through SFT could significantly degrade LLM in general-purpose language understanding tasks, hindering its practical deployment as a reranking agent in real-world scenarios.
+General rerankers achieve strong performance on listwise ranking tasks after SFT, but often lose their ability to follow general instructions.
+As shown in Figure 1, models like LLaMA may output incomplete or irrelevant results, such as numeric lists, when asked to write a travel blog.
+This suggests that SFT for ranking can reduce a model’s general text generation ability.
+In this study, we introduce a novel LLM reranking algorithm, namely RaCT, by implementing SFT through a chain-of-thought (CoT) prompt (see Fig.
+2) to guide sequential passage ranking without undermining the LLM’s general language understanding ability.
+To enhance the reasoning ability, we further incorporate CoT prompting into ranking preference optimization (RPO) Rafailov et al.
+(2024), proposing a new step-wise ranking preference learning framework RaCT.
+The framework utilizes overlapping ranking orders as the reward function, enabling efficient and flexible error correction.
 We summarize the contributions of this work as follows:
-  • We propose a novel chain-of-thoughts instruction (reranking) tuning that enables LLMs to rank passages based on relevance step by step.
-  • To our best knowledge, this is the first work to introduce the Supervised Fine-Tune–DPO pipeline in the context of text reranking.
-  • Using the same training data, our model outperforms both RankZephyr and the teacher model, RankGPT4, while retaining the general capabilities of LLMs.
+  • To the best of our knowledge, this is the first research study to investigate the trade-off between ranking utility and language modeling for the recent emerging LLM re-rankers.
+  • We propose a novel chain-of-thoughts instruction (reranking) tuning algorithm – RaCT – that enables LLMs to rank passages based on relevance step by step, and RPO to enable further ranking preference optimization.
+  • Empirical evidence on three ranking benchmarks, TREC Deep Learning Tracks Craswell et al.
+(2021a, 2020), BEIR Thakur et al.
+(2021), and BRIGHT Su et al.
+(2025), and the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al.
+(2020) suggests that our approach achieves state-of-the-art LLM re-ranking performance with the preservation of intrinsic language modeling capabilities.
 {mdframed}
@@
 Only respond with each step and the final answer, ensuring each passage is included once and ranked in descending relevance.
-Figure 2: ChainRank Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively.
+Figure 2: RaCT Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively.
 The prompt ensures step-by-step selection, removal, and ordering of passages, with an example illustrating the expected output format.
@@
 $Left:$ Stage 1 aims to teach the student model to perform text reranking.
-$Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform Chain DPO training.
-2.1 CoT Reranking Prompt Design
-Recent advancements in listwise text reranking, such as RankVicuna Pradeep et al.
-(2023a), RankZephyr Pradeep et al.
-(2023b), and related methods Liu et al.
-(2024), build on zero-shot techniques Ma et al.
-(2023); Sun et al.
-(2023), using prompt templates to reorder documents for metrics like nDCG Järvelin and Kekäläinen (2002).
-Given a query $q$ and documents ${d_{1},d_{2},...,d_{n}}$, the goal is to return a reordered list (e.g., $[1]>[2]>[3]...>[20]$).
-Our ChainRank strategy frames listwise reranking as a chain-of-thought (CoT) reasoning task, selecting the most relevant document iteratively until all are ranked.
-The CoT prompt for LLaMA is shown in Figure 2.
-2.2 Training Dataset
-We train ChainRank model using 35k GPT-3.5 and 5k GPT-4 labeled instances from Pradeep et al.
-(2023b), derived from randomly selected MS MARCO v1 queries.
-Pyserini Lin et al.
-(2021) retrieved 20 BM25 candidates per query, which RankGPT3.5 and RankGPT4 ordered as teacher models.
-Malformed generations, such as missing or duplicate identifiers, were removed to improve data quality.
-We randomly selected 90% of the query-document pairs for Stage 1 SFT training and 10% for Stage 2 DPO training, denoted as SFT data and DPO data.
-2.3 Training Stage 1: Supervised Fine-Tune
-As shown in Figure 3, during the Supervised Fine-Tuning (SFT) phase, we use SFT data, maintaining a zero-shot setup since RankGPT and ChainRank do not rely on human-labeled data.
-Before fine-tuning, the original LLaMA3-8B-Instruct model Dubey et al.
-(2024) failed to produce meaningful CoT reranking, simply replicating the example format without relevance-based results, highlighting the need for fine-tuning.
-The model weights are publicly available on HuggingFace.
+$Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform RPO training.
+2.1 CoT Reranking Prompt and Dataset
+RaCT formulates listwise reranking as a chain-of-thought (CoT) reasoning task, where the model iteratively selects the most relevant passage until all are ranked.
+The prompt (Figure 2) guides LLaMA to output ranked indices given a query and a set of passages.
+This builds on prior zero-shot reranking methods such as RankVicuna and RankZephyr Pradeep et al.
+(2023a, b); Liu et al.
+(2024), which use template-based prompts to optimize ranking metrics like nDCG Järvelin and Kekäläinen (2002).
+We train RaCT on 40k examples from Pradeep et al.
+(2023b), with 20 BM25-retrieved passages per query labeled by RankGPT3.5 or RankGPT4.
+The dataset includes three prompt formats: standard RankGPT-style prompts,RaCT CoT prompts, and RaCT CoT prompts where the model outputs the ranked list directly without explanation.
+We split 90% for supervised CoT tuning and 10% for RPO.
+Denoted as CoT tuning data and RPO data.
+The dataset is publicly available.
+[^1] https://huggingface.co/datasets/castorini/rank_zephyr_training_data
+2.2 RaCT: Two-Stage Training
+Training Stage 1: CoT Tuning As shown in Figure 3, during the Stage 1, we use CoT Tuning data with mixed prompt templates, maintaining a zero-shot setup since RankGPT and RaCT do not rely on human-labeled data.
+Before fine-tuning, the LLaMA3.1-8B-Instruct model Grattafiori et al.
+(2024) failed to perform CoT reranking, merely copying the example format without relevance.
+This highlights the need for fine-tuning.
+Model weights are available on Hugging Face.
 [^2] https://huggingface.co/meta-LLaMA/Meta-LLaMA-3-8B-Instruct
-We fully fine-tune the 8B parameter LLaMA-3 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format.
-Training on four NVIDIA A100 80GB GPUs took approximately 39 hours.
-2.4 Training Stage 2: Chain DPO
-After Stage 1, the ChainRank-SFT model generates three ranking predictions ($y$) on prompts ($x$) from DPO training data.
+We fully fine-tune the 8B parameter LLaMA3.1 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format.
+Training on four NVIDIA A100 80GB GPUs took approximately 20 hours.
+We use Maximum Likelihood Estimation (MLE) as our objective function, optimizing the model to generate the most probable ranking sequence given the input passages and queries.
+Training Stage 2: RPO After Stage 1, the RaCT model generates three ranking predictions ($y$) on prompts ($x$) from RPO data.
 Predictions are evaluated by overlapping ranking orders with ground-truth labels, creating a preference dataset with prompts (x), chosen steps ($s_{w}$), rejected steps ($s_{l}$), and overlapping steps ($s_{o}$).
-Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y={s_{1},s_{2},\dots,s_{n}}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$.
+Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y=\{s_{1},s_{2},\dots,s_{n}\}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$.
 Overlapping steps ($s_{o}$) are tracked until a divergence is found, after which only the initial contiguous overlaps are included in $s_{o}$, and later steps are categorized as chosen ($s_{w}$) or rejected ($s_{l}$).
@@
 This improves CoT reasoning’s sequential decision-making and model consistency.
-The Chain DPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately eight hours.
+The RPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately 6 hours.
 3 Experiments
-3.1 Research Questions
-Before conducting the experiments, we formulate several research questions to ensure that ChainRank could effectively address the challenges:
--RQ1: Does the CoT instruction tuning improve the step-by-step relevance ranking of passages compared to traditional ranking methods?
--RQ2: How does our model’s performance on text reranking tasks compare with existing models like RankZephyr and RankGPT4 across various datasets and task settings?
--RQ3: Does introducing DPO improve the performance and robustness of our model?
 3.1 Evaluation Benchmarks
+We evaluate ranking performance on standard academic benchmarks, including TREC DL19 and DL20 Craswell et al.
+(2020, 2021a), each with 100 BM25-retrieved passages per query and human relevance labels.
+To assess generalization, we include additional cross-domain benchmarks: BEIR Thakur et al.
+(2021), TREC DL21/DL22 Craswell et al.
+(2021b, 2022), and BRIGHT Su et al.
+(2025).
+Reranking is conducted using a sliding window strategy Sun et al.
+(2023), details explained in Appendix C.
+To verify general-purpose capabilities, we also evaluate on MMLU Hendrycks et al.
+(2020), testing 57 subject areas for text understanding post fine-tuning.
+3.2 Results
+Table 1 compares model performance on TREC and MMLU.
+Our 8B RaCT model outperforms all baselines, including the larger, closed-source RankGPT4333RankGPT is the only closed-source and largest model in the table., despite using only 90% of the training data.
+It also surpasses LLaMA3.1 + SFT, confirming the benefit of our CoT reranking prompt.
+In the RPO stage, RaCT achieves further gains with just one epoch and shows strong robustness: performance drops only 0.004 under tighter sliding windows, compared to over 0.02 for others.
+We also applied LoRA to fine-tune only the linear layers in Stage 1, reducing training time by 4 hours while achieving a competitive 0.747 on DL19.
+This surpasses RankZephyr and RankGPT, showing that a lightweight fine-tuning approach can still achieve strong ranking performance with our CoT prompting strategy.
+Models MSv1 MSv2 MMLU DL19 DL20 DL21 DL22 AVG.
+RankGPT4 0.750 0.704 0.684 0.509 0.864 BM25 0.506 0.480 0.446 0.269 N/A Gemma-7B 0.533 0.530 0.446 0.269 0.649 LLaMA3.1 0.669 0.641 0.639 0.429 0.720 DeepSeek-R1-8B 0.565 0.503 0.558 0.380 0.530 RankZephyr 0.742 0.709 0.600 0.408 0.000 RankVicuna 0.668 0.655 0.608 0.421 0.373 LLaMA3.1 + SFT 0.707 0.635 0.660 0.441 0.628 RaCT (Ours) 0.758 0.720 0.706 0.532 0.720
+Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks.
+All the reranking tasks are based on BM25 retrieval results.
+The bold values highlight the best performance.
+Model BEIR BRIGHT
+BM25 0.426 0.137
+RankZephyr 0.422 0.130
+RaCT 0.545 0.178
+Table 2: Average performance comparison on BEIR and BRIGHT benchmarks (nDCG@10), see detailed scores in Appendix F and G
+Table 2 presents the performance comparison of RaCT against BM25 and RankZephyr across various BEIR and BRIGHt benchmark datasets.
+Both LLM models are reranking based on BM25 retrieval results.
+The results show that RankZephyr, despite being a fine-tuned model, receives a lower average score than the traditional retriever BM25, indicating that its reranking effectiveness does not always surpass basic retrieval methods.
+Combined with the results from DL21 and DL22, these findings suggest that while traditional fine-tuning methods improve performance on in-scope datasets, they do not necessarily generalize well to out-of-scope datasets.
+In contrast, RaCT demonstrates stronger generalization across different retrieval tasks, reinforcing the effectiveness of incorporating CoT prompting and RPO in improving reranking capabilities beyond the training distribution.
+On MMLU, RaCT matches LLaMA3.1’s performance, while LLaMA3.1 w/o CoT shows a slight drop and RankVicuna performs poorly.
+RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU.
+Although our model was trained using a mixed Chain-of-Thought (CoT) reranking prompt, we use the same concise prompt as RankGPT during inference and directly output the final ranked list.
+This ensures that inference time remains unchanged.
+3.3 Ablation Study
+Setting nDCG@10 (DL22)
+[…]
+– CoT 0.407
+– RPO 0.504
+– CoT&RPO 0.429
+[…]
+Table 3: Ablation study with LLaMA3.1-8B as the base model. Each row removes or modifies components from the full RaCT setup and reports performance on TREC DL22 (nDCG@10).
+Table 3 presents an ablation study based on LLaMA3.1-8B.
+Removing CoT or RPO from the RaCTBM25 baseline leads to performance drops, with CoT having a larger impact (0.407 vs.
+0.504).
+Removing both results in 0.429, which is lower than using CoT alone but higher than using RPO alone, indicating that both components contribute and their combination is most effective.
+Using SPLADE++ED as the retriever further boosts performance to 0.683, showing the reranker’s compatibility with stronger retrieval systems.
+4 Conclusion and Future Work
+In this paper, we introduced RaCT, a zero-shot listwise reranker based on LLaMA3.1, and showed the effectiveness of CoT prompting and CoT–RPO training.
+Our model outperforms both open- and closed-source baselines, including RankZephyr and RankGPT4.
+For future work, we plan to adapt RaCT to other architectures (e.g., Mistral, Zephyr, QWEN) and train on more diverse, higher-quality datasets to enhance robustness and performance.
+5 Limitations
+While our model is based on LLaMA 3.1, which supports a 128k context window—significantly longer than RankVicuna (4k) and RankZephyr (8k)—we observe that using extremely long contexts by including all passages at once can negatively impact reranking performance.
+In some cases, this leads to worse results compared to a sliding window approach.
+This suggests that simply leveraging longer context lengths does not always yield better rankings, possibly due to attention dilution or ineffective utilization of distant context.
+In future work, we aim to explore strategies to better manage long-context inputs, such as adaptive passage selection or hierarchical attention mechanisms, to fully exploit the model’s context capacity without compromising ranking quality
+Appendix A Baseline Selection
+To illustrate the efficacy of RaCT, we select several baseline models, including the unsupervised retriever model BM25 Robertson and Zaragoza (2009); zero-shot prompt-based LLMs: RankGPT4444gpt-4-0613, Gemma-7B Team et al.
+(2024), LLaMA3.1-8B-Instruct Grattafiori et al.
+(2024) and DeepSeek-R1-Distill-Llama-8B DeepSeek-AI et al.
+(2025); fine-tuned LLM-based models (RankVicuna, RankZephyr).
+To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankZephyr but use our base model LLaMA3.1-8B-Instruct, denoted as LLaMA3.1 + SFT.
+Appendix B Evaluation Benchmarks (Expanded)
 We evaluate ranking capabilities using TREC DL19 Craswell et al.
 (2020) and DL20 Craswell et al.
-(2021) Tracks, derived from MS MARCO V1 Bajaj et al.
+(2021a) Tracks, derived from MS MARCO V1 Bajaj et al.
 (2016), with human-annotated relevance labels.
-TREC DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 candidate passages retrieved by BM25.
-Additionally, we include cross-domain datasets from BEIR Thakur et al.
-(2021), not seen by the fine-tuned LLMs, for broader analysis.
-Using a sliding window strategy with a size of 20 and a stride of 10, we rerank 100 passages per query.
-We use the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al.
-(2020) to evaluate whether our model retains its text understanding abilities after fine-tuning for ranking tasks.
-MMLU tests proficiency across 57 diverse subjects, from elementary to advanced professional topics.
-We compare our model’s performance on MMLU with baseline LLMs to ensure ranking fine-tuning does not compromise general-purpose language understanding and generation capabilities, maintaining its versatility in broader tasks.
-3.3 Results
-Table 1 compares model performance on TREC, BEIR, and MMLU benchmarks.
-Our 8B ChainRank model outperforms baselines, including RankGPT4  [^2] RankGPT is the only closed-source and largest model in the table., even with 90% of the training data, and surpasses RankLLaMA3 under identical training conditions, showcasing the effectiveness of our CoT reranking prompt.
-In the DPO stage, ChainRank-DPO improves across all datasets with just one training epoch, enhancing robustness to passage variations (Appendix B).
-On MMLU, ChainRank matches LLaMA3’s performance, while RankLLaMA3 shows a slight drop and RankVicuna performs poorly.
-RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU.
-Additional examples are in Appendix D.
-We present a figure to illustrate the trade-off between performance and inference cost when generating ranking orders at varying step intervals.
-Notably, our method achieves comparable latency to RankLLaMA3 (26.5 vs.
-25.0 seconds).
-By utilizing parallel distributed evaluation, we further reduce inference latency to 6.73 seconds on 4 A100 GPUs, demonstrating the efficiency of our approach.
-Models TREC BEIR MMLU DL19 DL20 NFC COVID FIQA AVG.
-RankGPT4 0.746 0.708 0.406 0.749 0.333 0.864 BM25 0.506 0.480 0.325 0.595 0.236 N/A Contriever 0.616 0.599 0.328 0.596 0.329 N/A Gemma-7B 0.533 0.530 0.338 0.573 0.236 0.649 LLaMA3 0.641 0.621 0.269 0.628 0.216 0.662 RankZephyr 0.742 0.709 0.331 0.592 0.230 0.000 RankVicuna 0.668 0.655 0.338 0.592 0.236 0.373 RankLLaMA3 0.730 0.635 0.335 0.672 0.330 0.628 ChainRank-SFT 0.752 0.714 0.353 0.766 0.342 0.663 ChainRank-DPO 0.755 0.717 0.358 0.772 0.342 0.663
-Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks.
-All the reranking tasks are based on BM25 retrieval results.
-The bold values in the table highlight the best performance across the respective benchmarks.
-Figure 4: FLOPs and nDCG@10 performance across different step intervals.
-The blue line represents the inference cost (FLOPs), while the orange line shows the ranking performance (NDCG@10).
-As the step interval increases, FLOPs decrease significantly, with a slight drop in nDCG@10 scores.
-4 Conclusion and Future Work
-In this paper, we propose a novel approach, ChainRank, a novel zero-shot listwise text reranking model built on LLaMA3.
-We demonstrated the effectiveness of CoT reranking prompt and our SFT–Chain DPO pipeline.
-The results of our experiments show that our model achieves superior performance compared to many open-source and closed-source LLMs, such as RankZephyr and RankGPT4.
-In future work, we plan to explore the application of our method to other models, such as Mistral, Zephyr, and LLaMA3.1, to evaluate its generalizability and performance across different architectures.
-Additionally, our future research could incorporate higher-quality datasets with varying numbers of passages per instance to enhance the diversity and robustness of the training data.
-Expanding the dataset in this manner could lead to improved overall performance.
-5 Limitations
-While this study provides valuable insights, several limitations should be acknowledged.
-Firstly, the need to include long examples of CoT formatting (Step1: …) in our prompts leads to increased training and inference times.
-Secondly, although LLaMA3 has a larger context window (8192 tokens) compared to Vicuna and Zephyr (4096 tokens), we are still limited by the maximum number of passages in our training data, which is set to 20.
-As a result, we are restricted to ranking within a window size of 20 and require nine passes to complete the reranking of 100 passages.
-Appendix A Baseline Selection
-To illustrate the efficacy of ChainRank, we select several baseline models, including the existing retriever models, BM25 Robertson and Zaragoza (2009) and Contriever Izacard et al.
-(2021), zero-shot prompt-based LLMs, RankGPT4 and Gemma-7B Team et al.
-(2024), fine-tuned LLM-based models (RankVicuna, RankZephyr), as well as our base model LLaMA3-8B-Instruct.
-To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankVicuna but use our base model LLaMA3-8B-Instruct, denoted as RankLLaMA3.
-Appendix B Window Size
-Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 RankLLaMA3 0.567 0.663 0.666 ChainRank-SFT 0.697 0.737 0.755 ChainRank-DPO 0.671 0.755 0.758
+DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 BM25-retrieved candidate passages.
+For broader generalization analysis, we include additional datasets not seen during training: (1) BEIR Benchmarks Thakur et al.
+(2021), (2) TREC DL21 Craswell et al.
+(2021b) and DL22 Craswell et al.
+(2022) based on MS MARCO V2, and (3) BRIGHT: A Reasoning-Intensive Benchmark Su et al.
+(2025).
+Reranking uses a sliding window strategy Sun et al.
+(2023) with window size 20 and stride 10 to process long sequences.
+Rankings are computed independently for each segment and aggregated across the query.
+To assess whether ranking fine-tuning affects broader model capabilities, we test on MMLU Hendrycks et al.
+(2020), which covers 57 diverse academic and professional subjects.
+We compare performance before and after ranking fine-tuning to ensure the model retains general language understanding and generation abilities.
+Appendix C Sliding Window Strategy
+In the sliding window strategy commonly used in reranking models (e.g., RankZephyr, RankGPT, and our method), passages appearing in multiple overlapping windows are ranked multiple times based on the local context of each window.
+Since most reranking evaluations focus on nDCG@10, only the top 10 passages from the final ranking are considered.
+This means that passages appearing in the final window retain their latest assigned ranking, while earlier rankings from overlapping windows are not explicitly reconciled.
+This approach follows the standard strategy used in previous works and provides a computationally efficient way to handle long passage lists.
+Appendix D Base Model Selection
+Models DL19 DL20 DL21 DL22
+LLaMA3-8B 0.752 0.710 0.693 0.502
+LLaMA3.1-8B 0.753 0.710 0.698 0.504
+LLaMA3.2-3B 0.718 0.697 0.690 0.517
+LLaMA3.2-1B 0.713 0.682 0.673 0.465
+Table 4: Performance Comparison of different base models on TREC Benchmarks (nDCG@10).
+All the models are trained with identical RaCT strategy.
+The bold values highlight the best performance across the respective datasets.
+Appendix E Window Size
+Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3.1 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 LLaMA3.1 + SFT 0.567 0.663 0.666 RaCT w/o RPO 0.697 0.737 0.753 RaCT 0.702 0.754 0.758
 Table 5: Performance of different models with different window size and step size, BM25 performs all of the retrieval stage, and all of the evaluations are performed on DL19
-Appendix C Explanation of CoT strategies
-To evaluate the trade-off between performance and inference cost, we trained our model to generate ranking orders after every 1, 3, 5, and 7 steps.
-An additional instruction was introduced to guide the model in returning ranking orders at these specified intervals.
-For comparison, we also included the performance of RankLLaMA3 to highlight the effectiveness of our CoT strategies
-Appendix D Examples of general LLM tasks
+Appendix F BEIR Evaluation
+Datasets BM25 RankZephyr RaCT
+NFC 0.322 0.331 0.392
+FEVER 0.651 0.618 0.823
+COVID 0.595 0.592 0.847
+NEWS 0.395 0.396 0.471
+SciFact 0.679 0.679 0.790
+HotPotQA 0.633 0.669 0.762
+NQ 0.306 0.375 0.600
+Qoura 0.789 0.719 0.807
+FIQA 0.236 0.230 0.435
+CFEVER 0.165 0.160 0.237
+Arguana 0.397 0.295 0.310
+Touche 0.442 0.434 0.414
+SCIDOCS 0.149 0.148 0.207
+Signal 0.330 0.316 0.347
+DBPedia 0.318 0.395 0.459
+Robust04 0.407 0.406 0.492
+Avg. 0.426 0.422 0.545
+Table 6: Performance Comparison on BEIR Benchmarks (nDCG@10). The bold values highlight the best performance across the respective subsets.
+Appendix G BRIGHT Evaluation
+Datasets BM25 RankZephyr RaCT
+Bio. 0.182 0.219 0.246
+Earth. 0.279 0.237 0.331
+Econ. 0.164 0.144 0.185
+Psy. 0.134 0.103 0.228
+Rob. 0.109 0.076 0.164
+Stack. 0.163 0.137 0.157
+Sus. 0.161 0.166 0.205
+Pony 0.043 0.065 0.137
+Leet. 0.247 0.247 0.253
+AOPs 0.065 0.068 0.072
+TheoT 0.021 0.020 0.082
+TheoQ 0.073 0.073 0.081
+Avg. 0.137 0.130 0.178
+Table 7: Performance Comparison on BRIGHT Benchmarks (nDCG@10). The bold values highlight the best performance across the respective subsets.
+Appendix H Examples of General LLM Tasks
 To illustrate the degradation in general text understanding and generation capabilities of RankVicuna and Rankzephyr, we provide two examples below.
-Due to the length of ChainRank’s response, we have omitted the middle portion of the answer.
-D.1 Example 1
+Due to the length of RaCT’s response, we have omitted the middle portion of the answer.
+H.1 Example 1
 Prompt Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions.
@@
 RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [11] > [12] > [13]
-ChainRank Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me.
+RaCT Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me.
 I had dreamed of visiting Hawaii for years, and finally, I was here.
@@
 I hope my story has inspired you to plan your own trip to Hawaii and experience the warmth and beauty of this incredible destination.
-D.2 Example 2
+H.2 Example 2
 Prompt Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided?
@@
 RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [3] > [4] > [5] > [6]
-ChainRank I’d be happy to help you write a resignation letter to your current employer.
+RaCT I’d be happy to help you write a resignation letter to your current employer.
 Here’s a sample template you can use as a guide: Dear [Recipient’s Name], I am writing to inform you of my decision to resign from my position as [Your Position] at [Company Name], effective [Date of Last Day of Work].